# The Low-Temperature Phenomenology of Gap Inhomogeneity in the Cuprate Superconductors: High-Energy Granularity, Low-Energy Homogeneity, and Spectral Kinks

### v1

@@ IX Conclusion
Our results will hopefully spur a reexamination of the proper interpretation of kinks and the low-/high-energy dichotomy seen in many STS experiments on the cuprates. It has long been thought that the kink energy is the superconducting energy scale, while the spectral-gap energy is the pseudogap scale brought about by the presence of coexisting order. We find that both of these energy scales can have the same origin—a broadly distributed and spatially extended inhomogeneous $d$-wave order parameter. If, as we strongly suspect, both of these energy scales in the cuprates originate from the same source, then many questions naturally arise for future experimental and theoretical work in this area centered on the interplay between the inhomogeneous $d$-wave superconductivity in these materials and the presence of coexisting order. In particular, it is worth asking what role coexisting order plays in the cuprates when a surprisingly large extent of the experimental phenomenology at low temperatures is adequately accounted for by gap inhomogeneity. More importantly, some of the discrepancies between our theoretical mean-field models and experiment are surely due to beyond-mean-field effects (e.g., correlation effects, superconducting fluctuations) or coexisting order not taken into account in our simulations. Our work provides a building block with which one can tease out the beyond-mean-field ingredients needed to fully charactize cuprate phenomenology.
@@ Acknowledgements.
It is a pleasure to thank S.-D. Chen, A. Kreisel, M. Pal, and especially W. A. Atkinson, P. J. Hirschfeld, and the late J. Zaanen for very inspiring discussions, collaborations, and encouragement, and J. C. S. Davis for graciously giving permission to reprint Fig. 25a. M.A.S. acknowledges the hospitality of the Stanford Institute for Materials and Energy Sciences, where this work was initiated.

### v2

@@ IX Conclusion
Our results will hopefully spur a reexamination of the proper interpretation of kinks and the low-energy–high-energy dichotomy seen in many STS experiments on the cuprates. It has long been thought that the kink energy is the superconducting energy scale, while the spectral-gap energy is the pseudogap scale brought about by the presence of coexisting order. We find that both of these energy scales can have the same origin—a broadly distributed and spatially extended inhomogeneous $d$-wave order parameter. If, as we strongly suspect, both of these energy scales in the cuprates originate from the same source, then many questions naturally arise for future experimental and theoretical work in this area centered on the interplay between the inhomogeneous $d$-wave superconductivity in these materials and the presence of coexisting order. In particular, it is worth asking what role coexisting order plays in the cuprates when a surprisingly large extent of the experimental phenomenology at low temperatures is adequately accounted for by gap inhomogeneity. More importantly, some of the discrepancies between our theoretical mean-field models and experiment are surely due to beyond-mean-field effects (e.g., correlation effects, superconducting fluctuations) or coexisting order not taken into account in our simulations. Our work provides a building block with which one can tease out the beyond-mean-field ingredients needed to fully characterize cuprate phenomenology.
@@ Acknowledgements.
It is a pleasure to thank S.-D. Chen, A. Kreisel, M. Pal, and especially W. A. Atkinson, P. J. Hirschfeld, and the late J. Zaanen for very inspiring discussions, collaborations, and encouragement, and J. C. S. Davis for graciously giving permission to reprint Fig. 24a. M.A.S. acknowledges the hospitality of the Stanford Institute for Materials and Energy Sciences, where this work was initiated, and UFIT Research Computing for providing computational resources and support that have contributed to the research results reported in this publication.
X Data Availability
The data that support the findings of this article are openly available [72].
Appendix A Effects of Adjusting the Width of the Order-Parameter Distribution
Figure 26: Plots of the binned average local density of states for $\overline{\Delta}=0.16$ and $l=8$, and different values of the width parameter $w$. Each LDOS spectrum is binned according to its spectral gap (with 10 bins in all), and all spectra in a given bin are then averaged over.
Figure 27: Plots of the spectral gap $\Delta_{S}$, the low-energy gap $\Delta_{L}$, and the kink energy $\Delta_{K}$ of the bin-averaged spectra for $\overline{\Delta}=0.16$, $l=8$ as a function of $w$. The open markers indicate values of the relevant quantities corresponding to each bin-averaged spectrum, while the filled markers show the average of the said quantity over all spectra. The size of each marker is proportional to the number of spectra within the bin.
Figure 28: Plots of the two-dimensional histograms for $\Delta_{T}$ and $\Delta_{S}$ ((a) to (d)); $\Delta_{T}$ and $\Delta_{L}$ ((e) to (h)); and $\Delta_{S}$ and $\Delta_{L}$ ((i) to (l)) for $\overline{\Delta}=0.16$ and $l=8$, and different values of the width parameter $w$. The same inhomogeneity configurations are used for different values of $w$; the widths of these distributions are rescaled by $w$.
In this section, we clarify the role of the width of the order-parameter distribution on the various observables of interest to us. To this end, we make use of the $\overline{\Delta}=0.16$, $l=8$ model (with $l\approx\xi$) first shown in Sec. IV. We use the same inhomogeneity configurations as presented in that section, but this time, we adjust the width of the distribution by a scaling factor $w$ so that the resulting width is $2w\overline{\Delta}$; the mean is fixed at $\overline{\Delta}$. (For instance, if $w=0.50$, then the values that $\Delta_{D,i}$ can take lie in the range $0.5\overline{\Delta},1.5\overline{\Delta}]$.) As we have already considered the maximally inhomogeneous cases in the paper ($w=1$), we restrict ourselves to $w<1$ and consider the following values of $w$: $0.25$, $0.50$, and $0.75$. (Cases where $w>1$ are indeed possible, but these necessarily feature order-parameter sign reversals, which introduce attendant new physics of their own, and we neglect these in our present paper.)
In Fig. 26, we plot the bin-averaged LDOS for different values of $w$. It can be seen that in the narrowest case ($w=0.25$), the various spectra are very homogeneous and exhibit only minimal variations; the inhomogeneity only becomes apparent at energies near $\omega\approx\overline{\Delta}$, and even then, the resulting gaps vary only very mildly. This is very similar to the results we obtained for the maximally inhomogeneous small-patch model ($l<\xi$, $w=1$) we studied in detail in Sec. III. As $w$ is increased, the kink energy becomes smaller, while the distribution of gaps becomes broader. One can see that regardless of the value of $w$ used, the kink energy tends to coincide with the smallest gaps in the system. (These observations are more succinctly plotted in Fig. 27, which shows bin-averaged quantities such as the kink energy and the spectral gap as a function of $w$.) In the maximally inhomogeneous ($w=1$) case, the smallest gaps are found in patches where $\Delta_{T}\approx 0$; in the non-maximally inhomogeneous cases, these smallest gaps are instead found in regions where $\Delta_{T}$ is nonzero. This makes it clear that the peculiarities of the maximally inhomogeneous model, with regions where the superconducting order parameter vanishes, are not essential to the physics we have highlighted in this paper. All that is needed is that the distribution of the order parameter be sufficiently broad to see the effects of interest to experimental studies of the cuprates.
Fig. 28 shows two-dimensional histograms featuring pairs of order-parameter measures for varying $w$. A noteworthy feature of the plots for all three pairs of measures is that $w<1$ histograms look simply like “squished” versions of the $w=1$ plots. This is most evident for the histograms for $\Delta_{T}$-$\Delta_{S}$: the overall profiles of the joint distributions look similar across $w$. It is interesting to compare the $w=0.25$ case (Fig. 28a) with the $l<\xi$ case described in Sec. III (Fig. 5a): the $\Delta_{S}$ distribution appears similar, even though the $\Delta_{T}$ distribution is much narrower for $w=0.25$ and $l=8$ than it is for $w=1$ and $l=1$. These results highlight the role that the inhomogeneity length scale plays in the resulting variation in the spectral gap $\Delta_{S}$: inhomogeneity with small length scales effectively results in a homogeneous quasiparticle response even when the order parameter is broadly distributed. Put differently: for a fixed width of the spectral-gap distribution, the width of the underlying order-parameter distribution has to be greater (smaller) if a smaller (larger) $l$ is used.
These results allow us to interpret previous findings in a new light. The “nematic glass” scenario studied by Lee et al. [37] features spatially extended order-parameter inhomogeneity with $l\approx\xi$, but with a low-energy–high-energy dichotomy which does not feature kinks at a parametrically lower energy than than the average spectral gap. Their results actually resemble ours for very narrow distributions of spatially extended order-parameter inhomogeneity (e.g., Fig. 26a) where the energy at which homogeneity is lost coincides with the (very narrowly distributed) spectral gap. Because the inhomogeneous order parameter in Lee et al. is obtained by self-consistency with a spatially uniform pairing interaction, it might be the case that the resulting distribution is narrow, hence the absence of kinks in their spectra.